Light is traveling from air to an ‘unknown material’. The angle of incidence is measured at 32.0o, and the angle of refraction is measured at 27.5o. What is the speed of light in this ‘unknown material’. 



1
Expert's answer
2020-12-31T09:28:25-0500

"\frac{\sin\alpha_1}{\sin\alpha_2}=\frac{v_1}{v_2}\to v_2=v_1\cdot\frac{\sin\alpha_2}{\sin\alpha_1}=3\cdot10^8\cdot\frac{\sin27.5°}{\sin32°}=261406960(m/s)"